# SURE-VQA: Systematic Understanding of Robustness Evaluation in Medical VQA Tasks

### v1

@@ A.1 Failures of traditional metrics
Examples of failures of traditional token-matching metrics are shown in Figure 7
@@ A.2 Prompts for Evaluation
Listing 1: Mistral Prompt for Evaluating Open-Ended Questions
<s>[INST] You are a helpful evaluator to evaluate answers to questions about biomedical images.
@@ 5. Five Stars: "Perfect Match"
[/INST]
</s>
Listing 2: Mistral Prompt for Evaluating Closed-Ended Questions
<s>[INST] You are a helpful evaluator to evaluate answers to questions about biomedical images.
@@
The questions are all close ended, therefore the answer is either correct or false, there are no states in between.
Where the stars have the following meaning:
@@ 0. Zero Star: "Incorrect"
1. One Star: "Perfect Match"
- The answer exactly matches the ground truth with no discrepancies.
@@ 1. One Star: "Perfect Match"
[/INST]
</s>
Listing 3: Mistral Prompt for Evaluating Closed-Ended Multilabel Questions
<s>[INST] You are a helpful evaluator to evaluate answers to questions about biomedical images.
@@ 0.5 Star: "Partially Correct"
- The answer partially matches the ground truth, but contains some inaccuracies.
- Demonstrates a partial understanding of the question, providing a partially correct response.
@@
1 Star: "Perfect Match"
- The answer exactly matches the ground truth with no discrepancies.
@@ Appendix B Human Rater Study Details
B.1 Detailed Results of the Human Rater Study
The following figures show detailed results of the human rater study. Figure 8, Figure 9, and Figure 10 show scatter plots with the correlation between the human ratings and the other metrics. Figure 11 shows detailed correlation results, including the correlation between Mistral and the other metrics.
(a) SLAKE Human - Mistral Correlation
@@ B.1 Detailed Results of the Human Rater Study
(a) OVQA Human - Mistral Correlation
@@
(a) MIMIC Human - Mistral Correlation
@@
Figure 11: Extended Results of the Human Rater Study. Human interrater correlation is calculated between two human raters. Shown are the Kendall and Spearman correlation between the human rating and all traditional metrics as well as the correlation between Mistral and the traditional metrics.
@@ Appendix C Robustness Study Details
C.1 Hyperparameter Search
We performed several hyperparameter sweeps for each dataset and PEFT method in order to find suitable setups for the experiments in the PEFT robustness study. For the hyperparameter sweeps, we trained on the whole training set for each dataset and PEFT method and ran inference on the validation set. Training ran for 3 epochs and 3 seeds for each experiment.
C.1.1 Prompt Tuning
  • $\bullet$
    Number of tokens: $[40,60,80,100]$
  • $\bullet$
    Learning rate: $[3e-2,3e-1]$
Table 1: Hyperparameter sweep for prompt tuning on the SLAKE dataset. Selected hyperparameters for the final PEFT robustness study are highlighted. Mean and standard deviation are reported for three seeds.
# Tokens Learning Rate Closed-Ended (Mistral Accuracy) Open-Ended (Mistral Score)
40 3e-2 0.79 +/- 0.02 4.17 +/- 0.04
40 3e-1 0.8 +/- 0.02 4.19 +/- 0.05
60 3e-2 0.76 +/- 0.04 4.17 +/- 0.03
60 3e-1 0.81 +/- 0.01 4.17 +/- 0.05
80 3e-2 0.78 +/- 0.05 4.15 +/- 0.01
80 3e-1 0.82 +/- 0.01 4.18 +/- 0.02
100 3e-2 0.77 +/- 0.06 4.15 +/- 0.05
100 3e-1 0.81 +/- 0.0 4.18 +/- 0.03
Table 2: Hyperparameter sweep for prompt tuning on the OVQA dataset. Selected hyperparameters for the final PEFT robustness study are highlighted. Mean and standard deviation are reported for three seeds.
# Tokens Learning Rate Closed-Ended (Mistral Accuracy) Open-Ended (Mistral Score)
40 3e-2 0.85 +/- 0.0 2.96 +/- 0.06
40 3e-1 0.85 +/- 0.01 2.95 +/- 0.05
60 3e-2 0.85 +/- 0.01 2.98 +/- 0.04
60 3e-1 0.85 +/- 0.0 2.97 +/- 0.04
80 3e-2 0.81 +/- 0.04 2.96 +/- 0.04
80 3e-1 0.84 +/- 0.01 2.99 +/- 0.02
100 3e-2 0.83 +/- 0.02 2.99 +/- 0.04
100 3e-1 0.85 +/- 0.0 3.0 +/- 0.03
Table 3: Hyperparameter sweep for prompt tuning on the MIMIC dataset. Selected hyperparameters for the final PEFT robustness study are highlighted. Mean and standard deviation are reported for three seeds.
# Tokens Learning Rate Closed-Ended (Mistral Accuracy) Open-Ended (Mistral Score)
40 3e-2 0.67 +/- 0.02 3.17 +/- 0.03
40 3e-1 0.67 +/- 0.01 3.15 +/- 0.04
60 3e-2 0.68 +/- 0.01 3.14 +/- 0.01
60 3e-1 0.69 +/- 0.01 3.19 +/- 0.02
80 3e-2 0.66 +/- 0.05 3.17 +/- 0.01
80 3e-1 0.68 +/- 0.02 3.17 +/- 0.03
100 3e-2 0.68 +/- 0.02 3.19 +/- 0.03
100 3e-1 0.67 +/- 0.01 3.17 +/- 0.03
C.1.2 LoRA
For LoRA, we performed the following hyperparameter sweeps:
  • $\bullet$
    Rank: $[16,32,64,128,256]$
  • $\bullet$
    Learning rate: $[3e-5,3e-4]$
$\alpha$ is set to $2\times\text{Rank}$. The results for SLAKE can be found in Table 4, for OVQA in Table 5, and for MIMIC in Table 6. Note that some of the hyperparameter configurations led to instabilities during training loss, indicated by ”NaN”.
Table 4: Hyperparameter sweep for LoRA on the SLAKE dataset. Selected hyperparameters for the final PEFT robustness study are highlighted. Mean and standard deviation are reported for three seeds. Rows with ”NaN” showed instabilities in the loss during training.
Rank Learning Rate Closed-Ended (Mistral Accuracy) Open-Ended (Mistral Score)
16 3e-5 0.83 +/- 0.01 4.24 +/- 0.02
16 3e-4 0.82 +/- 0.01 4.23 +/- 0.05
32 3e-5 0.85 +/- 0.01 4.27 +/- 0.04
32 3e-4 0.73 +/- 0.07 4.2 +/- 0.03
64 3e-5 0.84 +/- 0.01 4.29 +/- 0.04
64 3e-4 0.52 +/- 0.07 3.01 +/- 1.28
128 3e-5 0.84 +/- 0.01 4.31 +/- 0.03
128 3e-4 NaN NaN
256 3e-5 0.83 +/- 0.01 4.28 +/- 0.01
256 3e-4 NaN NaN
Table 5: Hyperparameter sweep for LoRA on the OVQA dataset. Selected hyperparameters for the final PEFT robustness study are highlighted. Mean and standard deviation are reported for three seeds. Rows with ”NaN” showed instabilities in the loss during training.
Rank Learning Rate Closed-Ended (Mistral Accuracy) Open-Ended (Mistral Score)
16 3e-5 0.84 +/- 0.0 3.02 +/- 0.07
16 3e-4 0.83 +/- 0.02 3.08 +/- 0.03
32 3e-5 0.85 +/- 0.0 3.04 +/- 0.01
32 3e-4 0.82 +/- 0.01 2.99 +/- 0.04
64 3e-5 0.85 +/- 0.0 3.11 +/- 0.02
64 3e-4 0.65 +/- 0.0 2.04 +/- 0.1
128 3e-5 0.85 +/- 0.0 3.09 +/- 0.04
128 3e-4 NaN NaN
256 3e-5 0.85 +/- 0.0 3.1 +/- 0.03
256 3e-4 NaN NaN
Table 6: Hyperparameter sweep for LoRA on the MIMIC dataset. Selected hyperparameters for the final PEFT robustness study are highlighted. Mean and standard deviation are reported for three seeds. Rows with ”NaN” showed instabilities in the loss during training.
Rank Learning Rate Closed-Ended (Mistral Accuracy) Open-Ended (Mistral Score)
16 3e-5 0.7 +/- 0.01 3.31 +/- 0.01
16 3e-4 0.68 +/- 0.01 3.18 +/- 0.04
32 3e-5 0.71 +/- 0.0 3.33 +/- 0.02
32 3e-4 0.42 +/- 0.16 2.34 +/- 0.06
64 3e-5 0.71 +/- 0.01 3.33 +/- 0.03
64 3e-4 NaN NaN
128 3e-5 0.7 +/- 0.0 3.35 +/- 0.04
128 3e-4 NaN NaN
256 3e-5 NaN NaN
256 3e-4 NaN NaN
C.2 $\mathtt{(IA)^{3}}$
  • $\bullet$
    Learning rate: $[3e-3,3e-2,3e-1]$
Table 7: Hyperparameter sweep for $\mathtt{(IA)^{3}}$on the SLAKE dataset. Selected hyperparameters for the final PEFT robustness study are highlighted. Mean and standard deviation are reported for three seeds.
Learning Rate Closed-Ended (Mistral Accuracy) Open-Ended (Mistral Score)
lr3e-3 0.63 +/- 0.02 3.74 +/- 0.02
lr3e-2 0.83 +/- 0.01 4.28 +/- 0.02
lr3e-1 0.65 +/- 0.01 4.21 +/- 0.05
Table 8: Hyperparameter sweep for $\mathtt{(IA)^{3}}$on the OVQA dataset. Selected hyperparameters for the final PEFT robustness study are highlighted. Mean and standard deviation are reported for three seeds.
Learning Rate Closed-Ended (Mistral Accuracy) Open-Ended (Mistral Score)
lr3e-3 0.75 +/- 0.01 2.84 +/- 0.02
lr3e-2 0.84 +/- 0.0 3.08 +/- 0.01
lr3e-1 0.78 +/- 0.04 2.97 +/- 0.05
Table 9: Hyperparameter sweep for $\mathtt{(IA)^{3}}$on the MIMIC dataset. Selected hyperparameters for the final PEFT robustness study are highlighted. Mean and standard deviation are reported for three seeds.
Learning Rate Closed-Ended (Mistral Accuracy) Open-Ended (Mistral Score)
lr3e-3 0.53 +/- 0.0 2.86 +/- 0.01
lr3e-2 0.7 +/- 0.01 3.3 +/- 0.04
lr3e-1 0.61 +/- 0.05 3.06 +/- 0.04
C.3 Dataset Details
C.3.1 SLAKE
The SLAKE dataset Liu et al. (2021a) is a bilingual radiological VQA dataset, containing English and Chinese questions. We use the English subset of the SLAKE dataset. The dataset is composed of MRI, CT, and X-ray images. All images are 2D, so for the MRI and CT images, single slices are extracted. For each question, metadata information about the location, the modality, and the content is provided. Overall, the images are split into 5 different body locations, 11 different content types (question types), and the mentioned three modalities.
@@ C.3.1 SLAKE
C.3.2 OVQA
The OVQA dataset Huang et al. (2022) is an orthopedic VQA dataset, containing CT and X-Ray images. All images are 2D, so for the CT images, either a 3D rendering is shown as a 2D image or a single plane. For each question, metadata information is provided about the imaged organ (like the ”location” in the SLAKE dataset), and the question type (like the ”content” in SLAKE) is provided. The dataset contains 6 different question types and 4 different body parts.
@@ C.3.2 OVQA
C.3.3 MIMIC-CXR-VQA
The MIMIC-CXR-VQA dataset Bae et al. (2023) is a chest X-ray dataset, which is built based on the MIMIC-CXR dataset Johnson et al. (2019), the MIMIC-IV dataset Johnson et al. (2023), and the Chest ImaGenome dataset Wu et al. (2021). For each question, the semantic type is specified. Three different semantic types are specified, which are ”choose”, ”query”, and ”verify”. For ”choose”, the task is to choose between two options provided in the answer, but also both or none of the options can be correct. For ”query”, the task is to list all the categories that match the questions, e.g. all anatomical findings. Lastly, ”verify” are yes/no questions. All the questions can be answered based on a fixed set of classes, where the dataset overall contains 110 answer labels. The answers are given as a list of the correct classes. We preprocess the questions differently, based on their semantic type: For the ”choose” questions, whenever the list of answers contains both options, we change the answer to ”both”, and whenever the list of answers is empty, we change the answer to ”none”. For the ”query” questions, we concatenate the list of answers to one string, with the answer labels being comma-separated. For the ”verify” questions, we do not apply any specific preprocessing.
The information for the patient’s gender, ethnicity, and age are taken from the MIMIC-IV dataset. Whenever the metadata information of a subject ID is not unique, we set it to ”none”. In the respective shifts, we exclude questions where the corresponding metadata field is not known, which includes all fields with ”none”, and for the ethnicity shift also the value ”unknown/other”. The exact sizes of the dataset splits are listed in Table 12.
@@ C.3.3 MIMIC-CXR-VQA
C.3.4 Ratio of unique questions in the datasets

### v2

@@ A.1 Failures of traditional metrics
Examples of failures of traditional token-matching metrics are shown in Figure 9.
@@ A.2 Prompts for Evaluation
Listing 1: Gemma Prompt for Evaluating Open-Ended Questions
You are a helpful evaluator to evaluate answers to questions about biomedical images.
@@ 5. Five Stars: "Perfect Match"
Input:
Listing 2: Gemma Prompt for Evaluating Closed-Ended Questions
You are a helpful evaluator to evaluate answers to questions about biomedical images.
@@
The questions are all close-ended, so the answer is either correct or incorrect, but minor variations in phrasing or acceptable synonyms should still count as correct if the core meaning remains unchanged.
Evaluate whether the prediction (pred) accurately matches the ground truth (gt) based on meaning and relevance.
The stars have the following meaning:
@@ 0. Zero Star: "Incorrect"
- The predicted answer is incorrect.
- The main entity or concept from the ground truth is not correctly identified in the prediction
1. One Star: "Correct"
- The predicted answer is correct.
- The main entity or concept from the ground truth is correctly identified in the prediction
- the prediction provides the same information or identifies the same entity/concept as ground truth even if it includes additional, irrelevant details.
- Ensure that unrelated phrases or extra descriptions in the prediction do not distract from the evaluation of its correctness.
@@ 1. One Star: "Correct"
Input:
Listing 3: Gemma Prompt for Evaluating Closed-Ended Multilabel Questions
You are a helpful evaluator to evaluate answers to questions about biomedical images.
@@
The questions are all close-ended, so the answer is either correct or incorrect, but minor variations in phrasing or acceptable synonyms should still count as correct if the core meaning remains unchanged.
Evaluate whether the prediction (pred) accurately matches the ground truth (gt) based on meaning and relevance.
@@ 0 Star: "Incorrect"
- The predicted answer is incorrect.
- The main entity or concept from the ground truth is not correctly identified in the prediction.
@@ 0.5 Star: "Partially Correct"
- The predicted answer is partially correct.
- The main entity or concept from the ground truth is partially correctly identified in the prediction.
@@
1 Star: "Correct"
- The predicted answer is correct.
- The main entity or concept from the ground truth is correctly identified in the prediction
- The prediction provides the same information or identifies the same entity/concept as ground truth even if it includes additional, irrelevant details.
@@ 1 Star: "Correct"
- Ensure that unrelated phrases or extra descriptions in the prediction do not distract from the evaluation of its correctness.
@@ Appendix B Human Rater Study Details
B.1 Comparison of different LLMs
Figure 10 compares various LLMs and traditional metrics based on their correlation with human raters. On the SLAKE and OVQA datasets, the best-performing LLMs surpass traditional metrics in aligning with human judgments. However, on the MIMIC dataset, traditional metrics slightly outperform the LLMs. Interestingly, models specialized for the medical domain (Biomistral and Biomistral DARE) or designed as evaluators (Prometheus) do not demonstrate any advantage and even rank among the worst-performing models. On average, Gemma, Qwen, and Mistral v0.3 achieve the highest correlation with human judgments, with Gemma notably ranking in the top three across all datasets. Among the traditional metrics, the F1 score exhibits the best mean performance across datasets.
Figure 10: Comparison of various LLMs and Traditional Metrics in the Human Rater Study. Human interrater correlation is calculated between five human raters. Shown are the Kendall correlation between the human rating and LLM metrics as well as the traditional metrics. Gold, silver, and bronze medal show the best LLMs on the respective datasets. Left side shows the individual datasets and right side the mean over all datasets.
B.2 Qualitative Results of the Human Rater Study
Figure 11 shows qualitative results for the human rater study. Thereby, 11(a) shows failures of the traditional metric on the OVQA dataset, where the LLM metrics clearly outperform the traditional metrics in terms of correlation to humans. In contrast 11(b) show failures of Gemma on the MIMIC dataset, where the traditional metrics show a slightly higher correlation with human judgment.
(a) Failures of traditional metrics on the OVQA dataset
(b) Failures of Gemma on the MIMIC dataset.
Figure 11: Qualitative Results of the Human Rater Study. For each sample, the question, ground truth, prediction, and expected score by human ratings are shown. On the bottom, for each automated metric, the absolute value is shown with an indication if it is high or low in the metrics range and an indication if the expectations from the human ratings are met. Gemma scores range from 1-5 and traditional metrics from 0-1.
B.3 Detailed Quantitative Results of the Human Rater Study
The following figures show detailed results of the human rater study. Figure 12, Figure 13, and Figure 14 show scatter plots with the correlation between the human ratings and the other metrics. Figure 15 shows detailed correlation results, including the correlation between Gemma and the other metrics.
(a) SLAKE Human - Gemma Correlation
@@ B.3 Detailed Quantitative Results of the Human Rater Study
(a) OVQA Human - Gemma Correlation
@@
(a) MIMIC Human - Gemma Correlation
@@
Figure 15: Extended Results of the Human Rater Study. Human interrater correlation is calculated between five human raters. Shown are the Kendall and Spearman correlation between the human rating and all metrics as well as the correlation between Gemma and the traditional metrics.
@@ Appendix C Robustness Study Details
C.1 Dataset Details
C.1.1 SLAKE
The SLAKE dataset (Liu et al. (2021)) is a bilingual radiological VQA dataset, containing English and Chinese questions. We use the English subset of the SLAKE dataset. The dataset is composed of MRI, CT, and X-ray images. All images are 2D, so for the MRI and CT images, single slices are extracted. For each question, metadata information about the location, the modality, and the content is provided. Overall, the images are split into 5 different body locations, 11 different content types (question types), and the mentioned three modalities.
@@ C.1.1 SLAKE
C.1.2 OVQA
The OVQA dataset (Huang et al. (2022)) is an orthopedic VQA dataset, containing CT and X-Ray images. All images are 2D, so for the CT images, either a 3D rendering is shown as a 2D image or a single plane. For each question, metadata information is provided about the imaged organ (like the "location" in the SLAKE dataset), and the question type (like the "content" in SLAKE). The dataset contains 6 different question types and 4 different body parts.
@@ C.1.2 OVQA
C.1.3 MIMIC-CXR-VQA
The MIMIC-CXR-VQA dataset (Bae et al. (2023)) is a chest X-ray dataset, which is built based on the MIMIC-CXR dataset (Johnson et al. (2019)), the MIMIC-IV dataset (Johnson et al. (2023)), and the Chest ImaGenome dataset (Wu et al. (2021)). For each question, the semantic type is specified. Three different semantic types are specified, which are "choose", "query", and "verify". For "choose", the task is to choose between two options provided in the answer, but also both or none of the options can be correct. For "query", the task is to list all the categories that match the questions, e.g. all anatomical findings. Lastly, "verify" are yes/no questions. All the questions can be answered based on a fixed set of classes, where the dataset overall contains 110 answer labels. The answers are given as a list of the correct classes. We preprocess the questions differently, based on their semantic type: For the "choose" questions, whenever the list of answers contains both options, we change the answer to "both", and whenever the list of answers is empty, we change the answer to "none". For the "query" questions, we concatenate the list of answers to one string, with the answer labels being comma-separated. For the "verify" questions, we do not apply any specific preprocessing.
The information for the patient’s gender, ethnicity, and age is taken from the MIMIC-IV dataset. Whenever the metadata information of a subject ID is not unique, we set it to "none". In the respective shifts, we exclude questions where the corresponding metadata field is not known, which includes all fields with "none", and for the ethnicity shift also the value "unknown/other". The exact sizes of the dataset splits are listed in Table 3.
@@ C.1.3 MIMIC-CXR-VQA
C.1.4 Ratio of unique questions in the datasets